# Can maiBERT Speak for Maithili?

### v2

@@ Abstract
Can maiBERT Speak for Maithili?
Sumit Yadav1, Raju Kumar Yadav1, Utsav Maskey2, Gautam Siddharth Kashyap2
Md Azizul Hoque2, Ganesh Gautam1 1IOE, Pulchowk Campus, Lalitpur, Nepal {076bct088.sumit, 076bct100.raju, ganesh}@pcampus.edu.np 2Macquarie University, Sydney, Australia utsav.maskey@mq.edu.au gautam.kashyap@hdr.mq.edu.au mdazizul.hoque@students.mq.edu.au
@@ 1 Introduction
Figure 1: Density-Based Map of Maithili Language
@@
To bridge this critical gap, we propose a novel approach that focuses on the development of a Maithili-specific language model using the BERT architecture. Our work begins with the construction of a high-quality Maithili corpus collected from diverse sources, including online articles, books, and conversational data. This raw corpus forms the basis for pre-training a Masked Language Model (MLM) tailored specifically for Maithili. Following pre-training, we fine-tune the model for a downstream NLP task—text classification. Our model, referred to as maiBERT, is then benchmarked against State-of-the-Art (SOTA) to evaluate performance. Moreover, to support the wider research and developer community, we have open-sourced the pre-trained maiBERT model on Hugging Face.
@@ 2 Dataset
2.1 Pre-training Corpus:
The raw corpus [^3] https://huggingface.co/datasets/rockerritesh/maithiliNewsData for MLM pre-training consists of 1,028,017 sentences, with an average sentence length of 16.88 words and an average word length of 4.44 characters. Figure 2 provides an overview of the Maithili text corpus, illustrating how characters, words, and key linguistic features are distributed across the dataset.
Figure 3: Word distribution patterns in Maithili corpus.
Classification Dataset: For downstream evaluation, we constructed a separate labeled dataset from multiple credible Maithili news sources, such as Esamaad [^5] https://esamaad.com/, Maithilijindabad [^6] https://maithilijindabaad.com/, and Maithili Purnajagran Prakash [^7] https://mppdainik.com/.
@@ 2.1 Pre-training Corpus:
3 Related Works
NLP has shown remarkable advancements in foundational tasks such as semantic similarity and text categorization, significantly benefiting high-resource languages but posing persistent challenges for low-resource languages like Maithili (Chandrasekaran and Mago, 2021; Zhang et al., 2015). While semantic similarity tasks often depend on pretrained models or lexical resources, distributional semantic models such as Latent Semantic Analysis (LSA) (Landauer et al., 1998), Word2Vec (Church, 2017), and GloVe (Schnabel et al., 2015) have proven effective in capturing word relationships through co-occurrence patterns. However, these models often suffer from a lack of interpretability and are limited by the quality and domain-specificity of the training corpus. Therefore, Hadj Taieb et al. (2015) proposed a feature-based measure of sentence semantic similarity using noun, verb, and common word similarities on the Microsoft Research Paraphrase Corpus [^8] https://www.microsoft.com/en-us/download/details.aspx?id=52398 (MRPC), achieving a 79% agreement with human perception; however, their approach is heavily dependent on handcrafted linguistic features, limiting generalizability across domains and languages. Colla et al. (2020) introduced two novel similarity metrics and sense-level embeddings using six different methods, showing improved granularity in meaning representation; nonetheless, their method suffers from high computational complexity and lacks adaptation to low-resource languages. Nguyen et al. (2019) proposed short-text semantic similarity models through pretrained embeddings and external knowledge sources, evaluating them on MRPC, STS2015999https://github.com/brmson/dataset-sts, and P4PIN (Álvarez-Carmona et al., 2018) with promising Pearson correlation scores of 84.17%, 83.73%, and 95.22%, respectively; still, the study does not investigate robustness across morphologically rich or low-resource languages. Hartmann et al. (2017) trained 31 word embedding models for Portuguese-Brazilian using Word2Vec, FastText [^10] https://huggingface.co/blog/fasttext, GloVe, and Wang2Vec [^11] https://github.com/wlin12/wang2vec and recommended task-specific evaluations for syntactic and semantic similarity; however, the findings are limited to Portuguese and cannot be directly extrapolated to under-resourced Indic languages. Alian and Awajan (2020) reviewed semantic similarity techniques in Arabic and English and concluded that feature-based models outperform others; nonetheless, their study lacks empirical validation on cross-lingual or low-resource language scenarios. Pandit et al. (2019) proposed path-based and distributional models for cross-lingual semantic similarity, identifying Word2Vec as superior; yet, their evaluation lacks coverage for languages like Maithili.
With the advent of deep learning architectures like Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), text categorization tasks have also significantly improved (D’silva and Sharma, 2022), though early studies did not focus on low-resource settings or multilingual adaptability. Therefore, to address the data scarcity in low-resource languages, transfer learning has emerged as a widely adopted strategy where pretrained models from high-resource languages are fine-tuned on target languages (Ali et al., 2022). Patil and Kolhe (2022) demonstrated this by adapting an English sentiment model for Marathi through parameter optimization, although their study lacks scalability for languages with significantly different syntax and morphology. Similarly, Nie et al. (2022) proposed a novel cross-lingual emotion categorization model for Swahili and Urdu, but the absence of domain-specific datasets restricts real-world applicability. Ramchandra et al. (2019) explored semantic similarity in Hindi using multilingual deep learning models such as BERT and LASER, highlighting their comparative effectiveness; however, their results do not consider dialectal or regional variations commonly found in Hindi’s linguistic relatives. Lalrempuii et al. (2021) developed a NER framework for Assamese using manually annotated data and neural networks, achieving reasonable accuracy; however, the manual annotation process is labor-intensive and limits scalability. Priyadarshi and Saha (2020) created a Maithili POS tagger using CRF and neural word embeddings, reaching 85.88% accuracy with a 52,190-word corpus, but the system’s performance may degrade for out-of-vocabulary or code-switched text. Mundotiya et al. (2022) developed annotated corpora and POS/chunking tools for Bhojpuri, Maithili, and Magahi, showing high accuracy for SAHBiLC and FineSAHBiLC models (e.g., 0.86%–0.95% for POS and chunking); still, their approach is limited by the static nature of the corpus and its lack of domain diversity. Note: Word vectorization techniques, in general, have proven to be transformational for low-resource languages, particularly since 2013, yet the challenge remains in crafting embeddings that reflect the linguistic and morphological intricacies of underrepresented languages without relying excessively on sister-language transfer models. While transfer learning and distributional models enable rapid development, they often inherit biases and linguistic mismatches from their source languages.
@@ 4 Methodology
where $\mathcal{M}\subseteq\{1,\ldots,n\}$ denotes the masked positions and $\mathbf{x}_{\mathcal{M}}$ represents the input sequence with masked tokens. We train a Mithila language Language model with text corpus sourced from MithilaTextCorpus-v1 (refer to Section 2.1), ensuring a representative distribution of regional grammar, morphology, and vocabulary richness. The corpus was tokenized using SentencePiece-BPE, a subword-level tokenizer based on Byte-Pair Encoding (BPE), yielding a vocabulary $\mathcal{V}$ of size 30,000, ensuring subword-level coverage of rare and compound words. Each training sample was tokenized into sequences of length $n\leq 512$, padded or truncated appropriately. The final vocabulary embedding matrix ${W}_{e}\in{R}^{|\mathcal{V}|\times d}={R}^{30000\times 768}$ was learned jointly with model parameters. Training was conducted over 1,000,000 steps with a batch size of 64 and a learning rate $\eta=5\times 10^{-5}$, using the Adam optimizer with bias correction and weight decay regularization. Table 2 shows the training progress across key epochs, demonstrating rapid convergence of both loss and perplexity metrics. Gradient updates followed the update rule according to Equation (3):
@@ 5.2 Baselines
Note: In the tables, Blue highlights the best performance, while Green indicates the second-best.
@@ 6.1 LLM Pre-training
6.2 Downstream Classification and Comparison with SOTA
Table 3 demonstrates the performance of maiBERT. It handles Maithili news classification much better compared to existing state-of-the-art multilingual and Indic-specific baseline models. This consistent outperformance across all metrics can be attributed to several architectural and corpus-specific factors unique to maiBERT. First, unlike generic multilingual models such as mBERT (Gonen et al., 2020) and MuRIL (Khanuja et al., 2021), which are trained across a wide range of languages with varying resource availability, maiBERT is explicitly fine-tuned on a richly preprocessed and linguistically diverse Maithili corpus. This specialization enables it to better capture syntactic, semantic, and morphological nuances specific to the Maithili language, such as complex verb conjugations, dialectal variations, and usage of regional idioms, which are often underrepresented in large-scale multilingual pretraining. Moreover, the customized tokenization strategy leveraging a modified BERT tokenizer that preserves Devanagari-specific features plays a crucial role in improving embedding granularity, thereby enhancing context representation and classification robustness. The strong annotation quality and class-balanced labeled dataset further reinforce the model’s learning capabilities, where high inter-annotator agreement (Cohen’s Kappa = 0.82) ensures low label noise and improved generalization. When compared to HindiBERT (Joshi, 2022) and NepBERTa (Timilsina et al., 2022)—both of which are tailored for Indic languages—maiBERT still maintains a competitive edge due to its exclusive focus on Maithili language resources and domain-specific data collected from authentic news platforms. The improvement is particularly notable, with maiBERT achieving 87.02% accuracy and 86.9% macro-F1, outperforming the closest competitor (HindiBERT (Joshi, 2022) at 86.89% accuracy and 86.3% F1) by a meaningful margin. This suggests that maiBERT exhibits strong transfer capabilities and inherent language adaptability without needing extensive task-specific tuning. In contrast, mBERT (Gonen et al., 2020) and MuRIL (Khanuja et al., 2021) perform considerably lower, highlighting the limitations of overly generalized multilingual representations when applied to underrepresented and morphologically rich languages such as Maithili. Additionally, the inclusion of localized syntactic cues and lexical patterns during pretraining equips maiBERT with superior contextual disambiguation abilities, which proves critical in multi-class classification tasks where inter-class boundaries may be subtle.
6.3 Computational Experiments

### v3

# MaiBERT: A Pre-training Corpus and Language Model for Low-Resourced Maithili Language

@@ Abstract
MaiBERT: A Pre-training Corpus and Language Model for Low-Resourced Maithili Language
Sumit Yadav1, Raju Kumar Yadav1, Utsav Maskey2, Gautam Siddharth Kashyap2,
Ganesh Gautam1, Usman Naseem2 1IOE, Pulchowk Campus, Lalitpur, Nepal {076bct088.sumit, 076bct100.raju, ganesh}@pcampus.edu.np 2Macquarie University, Sydney, Australia {utsav.maskey,usman.naseem}@mq.edu.au gautam.kashyap@hdr.mq.edu.au
@@ 1 Introduction
Figure 1: Density-Based Map of Maithili Language.
@@
To bridge this critical gap, we propose a Maithili-specific language model using the BERT architecture. Our work begins with the construction of a high-quality Maithili corpus collected from diverse sources, including online articles, books, and conversational data. This raw corpus forms the basis for pre-training a Masked Language Model (MLM) tailored specifically for Maithili. Following pre-training, we fine-tune the model for a downstream NLP task—text classification. Our model, referred to as maiBERT, is then benchmarked against State-of-the-Art (SOTA) to evaluate performance.
2 Related Works
NLP has shown remarkable advancements in foundational tasks such as semantic similarity and text categorization, significantly benefiting high-resource languages but posing persistent challenges for low-resource languages like Maithili (Chandrasekaran and Mago, 2021; Zhang et al., 2015). Therefore, to address the data scarcity in low-resource languages, transfer learning has emerged as a widely adopted strategy, where pretrained models from high-resource languages are fine-tuned on target languages (Ali et al., 2022). Patil and Kolhe (2022) demonstrated this by adapting an English sentiment model for Marathi through parameter optimization, although their study lacks scalability for languages with significantly different syntax and morphology. Similarly, Nie et al. (2022) proposed a novel cross-lingual emotion categorization model for Swahili and Urdu, but the absence of domain-specific datasets restricts real-world applicability. Ramchandra et al. (2019) explored semantic similarity in Hindi using multilingual deep learning models such as BERT and LASER, highlighting their comparative effectiveness; however, their results do not consider dialectal or regional variations commonly found in Hindi’s linguistic relatives. Lalrempuii et al. (2021) developed a NER framework for Assamese using manually annotated data and neural networks, achieving reasonable accuracy; however, the manual annotation process is labor-intensive and limits scalability. Priyadarshi and Saha (2020) created a Maithili POS tagger using CRF and neural word embeddings, reaching 85.88% accuracy with a 52,190-word corpus, but the system’s performance may degrade for out-of-vocabulary or code-switched text. Mundotiya et al. (2022) developed annotated corpora and POS/chunking tools for Bhojpuri, Maithili, and Magahi, showing high accuracy for SAHBiLC and FineSAHBiLC models (e.g., 0.86%–0.95% for POS and chunking); still, their approach is limited by the static nature of the corpus and its lack of domain diversity. Word vectorization techniques, in general, have proven to be transformational for low-resource languages, yet the challenge remains in crafting embeddings that reflect the linguistic and morphological intricacies of underrepresented languages without relying excessively on sister-language transfer models. While transfer learning and distributional models enable rapid development, they often inherit biases and linguistic mismatches from their source languages.
@@ 3 Dataset
3.1 Pre-training Corpus:
With the objective of training language models for Maithili, we gathered approximately 18 million words by combining and de-duplicating data from three sources: (1) Maithili Wikipedia articles [^3] https://mai.wikipedia.org ~880K words), (2) Digitized books [^4] https://archive.org/details/maithili-books/ obtained through OCR processing of 418 scanned documents sourced from arXiv and other archives (~10.4 million words), and (3) raw newspaper articles, such as Esamaad [^5] https://esamaad.com/, Maithilijindabad [^6] https://maithilijindabaad.com/, and Maithili Purnajagran Prakash [^7] https://mppdainik.com/ (~6.6 million words). The OCR extraction was performed using Tesseract Smith (2007) on Maithili literary works and historical texts. After preprocessing and merging, the final corpus comprises over 450MB of cleaned Maithili text, making it one of the largest publicly compiled datasets for this language.
The combined raw corpus [^8] https://huggingface.co/datasets/rockerritesh/maithiliNewsData for MLM pre-training consists of 1,028,017 sentences, with an average sentence length of 16.88 words and an average word length of 4.44 characters. Figure 2 provides an overview of the Maithili text corpus, illustrating how characters, words, and key linguistic features are distributed across the dataset.
Figure 3: Word distribution patterns in Maithili corpus.
Classification Dataset: For downstream evaluation, we constructed a separate labeled test set from the Maithili news sources, as mentioned in the pre-training corpus.
@@ 4 Methodology
where $\mathcal{M}\subseteq\{1,\ldots,n\}$ denotes the masked positions and $\mathbf{x}_{\mathcal{M}}$ represents the input sequence with masked tokens. We train a Mithila language Language model with text corpus sourced from MithilaTextCorpus-v1 (refer to Section 3.1), ensuring a representative distribution of regional grammar, morphology, and vocabulary richness. The corpus was tokenized using SentencePiece-BPE, a subword-level tokenizer based on Byte-Pair Encoding (BPE), yielding a vocabulary $\mathcal{V}$ of size 30,000, ensuring subword-level coverage of rare and compound words. Each training sample was tokenized into sequences of length $n\leq 512$, padded or truncated appropriately. The final vocabulary embedding matrix ${W}_{e}\in{R}^{|\mathcal{V}|\times d}={R}^{30000\times 768}$ was learned jointly with model parameters. Training was conducted over 1,000,000 steps with a batch size of 64 and a learning rate $\eta=5\times 10^{-5}$, using the Adam optimizer with bias correction and weight decay regularization. Gradient updates followed the update rule according to Equation (3):
@@ 6.1 LLM Pre-training
6.2 Downstream Classification and Comparison with SOTA
Table 3 demonstrates the performance of maiBERT. It handles Maithili news classification much better compared to existing state-of-the-art multilingual and Indic-specific baseline models. This consistent outperformance across all metrics can be attributed to several architectural and corpus-specific factors unique to maiBERT. First, unlike generic multilingual models such as mBERT (Gonen et al., 2020) and MuRIL (Khanuja et al., 2021), which are trained across a wide range of languages with varying resource availability, maiBERT is explicitly fine-tuned on a richly preprocessed and linguistically diverse Maithili corpus. This specialization enables it to better capture syntactic, semantic, and morphological nuances specific to the Maithili language, such as complex verb conjugations, dialectal variations, and usage of regional idioms, which are often underrepresented in large-scale multilingual pretraining. Moreover, the customized tokenization strategy leveraging a modified BERT tokenizer that preserves Devanagari-specific features plays a crucial role in improving embedding granularity, thereby enhancing context representation and classification robustness. The strong annotation quality and class-balanced labeled dataset further reinforce the model’s learning capabilities, where high inter-annotator agreement (Cohen’s Kappa = 0.82) ensures low label noise and improved generalization.
When compared to HindiBERT (Joshi, 2022) and NepBERTa (Timilsina et al., 2022)—both of which are tailored for Indic languages—maiBERT maintains a competitive edge due to its exclusive focus on Maithili language resources and domain-specific data collected from authentic news platforms. HindiBERT performs close to maiBERT (86.89% vs. 87.02%) on Maithili text, largely due to their shared Devanagari script, 60–70% lexical similarity, and standardized vocabulary in news domains. However, maiBERT outperforms HindiBERT by 5–7% in texts where Maithili-specific vocabulary and grammatical structures are more significant.
MaiBERT exhibits strong transfer capabilities and inherent language adaptability without needing extensive task-specific tuning. In contrast, mBERT (Gonen et al., 2020) and MuRIL (Khanuja et al., 2021) perform considerably lower, highlighting the limitations of overly generalized multilingual representations when applied to underrepresented and morphologically rich languages such as Maithili. Additionally, the inclusion of localized syntactic cues and lexical patterns during pretraining equips maiBERT with superior contextual disambiguation abilities, which proves critical in multi-class classification tasks where inter-class boundaries may be subtle.
6.3 Computational Experiments